Please answer Q2 (A and B) Will Rate Thank you!

A) Calculate index of refraction (n) of a medium where light can pass through at a speed of 1.5 x10^8 m/s?

B) What is the cut off wavelength (λc) for a photoelectric materials of work function 1.1 eV?
Answer #1

Refractive index

n=\frac{C}{v}    where C is velocity of light in vacuum and v is the velocity of light at medium

n=\frac{3*10^{8}}{1.5*10^{8}}=2

Refractive index =2

b)

Work function

W=\frac{hC}{\lambda _{0}}      where h is planks consant, C is velocity of light in vacuum, {\lambda _{0}} is cut of wavelength

{\lambda _{0}}=\frac{hC}W=\frac{6.626*10^{-34}*3*10^{8}}{1.1*1.602*10^{-19}}=1.127*10^{-6}m=1127nm